OpenStudy (anonymous):

Solve for x: 4|2x - 2| + 10 = 34 x = -4, x = 2 x = 4, x = -4 x = 2, x = -2 x = 4, x = -2

hartnn (hartnn):

start by subtracting 10 from both sides

hartnn (hartnn):

we are going to use this fact in this problem \(|a|= b \implies a=b \quad or \quad a =-b\)

OpenStudy (anonymous):

-6 |2x-2| = 34

OpenStudy (anonymous):

is it A?

hartnn (hartnn):

not actually \(4 |2x-2| +10-10 - 34-10 \\ 4|2x-2| = 24\) did u get this ?

OpenStudy (anonymous):

ohh i subtracted it wrong

hartnn (hartnn):

now divide both sides by 4 :)

OpenStudy (anonymous):

1|2x-2| = 6

hartnn (hartnn):

correct! so, when |a| = b, we get a=b or a =-b so here, for |2x-2| = 6 we get 2x-2 = 6 or 2x-2 = -6 did u get this ? can you solve these 2 equations ?

OpenStudy (anonymous):

so i solve 2x-2=6 and 2x-2=-6

OpenStudy (anonymous):

??

hartnn (hartnn):

yes, can u solve thosE? start by adding 2 on both sides

OpenStudy (anonymous):

so 4x-2=8 and 4x-2=-8

hartnn (hartnn):

the 2x would remain 2x only because 2x and 2 are not like terms

hartnn (hartnn):

2x -2 =6 so 2x-2+2 = 6+2 2x = 8 got this ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so what does x equal ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so its B?

hartnn (hartnn):

nopes for 2x-2 = -6 2x = -6+2 = -4 x= -2

hartnn (hartnn):

so its D

OpenStudy (anonymous):

ohhh ok thanks

hartnn (hartnn):

welcome ^_^